### v3

# On the sparsity of non-diagonalisable integer matrices and matrices with a given discriminant

Abstract.
We consider the set ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ of $n\times n$-matrices with integer elements of size at most $H$ and obtain upper bounds on the number of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$, for which the characteristic polynomial has a fixed discriminant $d$. When $d=0$, this corresponds to counting matrices with a repeated eigenvalue, and thus is related to counting non-diagonalisable matrices. For $d\neq 0$, this problem seems not to have been studied previously, while for $d=0$, both our approach and the final result improve on those of A. J. Hetzel, J. S. Liew and K. Morrison (2007).
Key words and phrases:
Integer matrix, non-diagonalisable matrix, characteristic polynomial, discriminant
2020 Mathematics Subject Classification:
11C20, 15B36, 15B52
1. Introduction
1.1. Set-up and motivation
For positive integers $n$ and $H$, we let ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ denote the set of all $n\times n$ matrices
$A=\left(a_{ij}\right)_{1\leq i,j\leq n}$
with integer entries of size $|a_{ij}|\leq H$. In particular, ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ is of cardinality ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)=\left(2H+1\right)^{n^{2}}$.
Various problems of arithmetic statistics with matrices from the set ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ have received a lot of attention see, for example, [1, 2, 8, 13, 18] for some recent results and further references. Here we consider a rather natural question of counting the number $N_{n}(H)$ of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ which are not diagonalisable.
We note that previously this question has been studied by Hetzel, Liew and Morrison [19], where in fact the number $R_{n}(H)$ of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with a repeated eigenvalue is estimated (observe that we trivially have $R_{n}(H)\geq N_{n}(H)$). Note that [19, Theorem 2.1] merely asserts $R_{n}(H)=o(H^{n^{2}})$ for a fixed $n$ and $H\to\infty$. However, analysing the proof of [19, Theorem 2.1], exploiting the vanishing of the discriminant of the characteristic polynomial of such matrices, one easily sees that it actually leads to the bound
(1.1) $R_{n}(H)=O\left(H^{n^{2}-1}\right),$
where hereafter the implied constants depend only on $n$.
As noted in Remark 2.2 below, conditionally on [18, Conjecture 1.1] on the upper bound on the number of matrices with a give characteristic polynomial, see (2.1) below, it is natural to expect that
(1.2) $R_{n}(H)\leq H^{n^{2}-2n+2+o(1)}.$
One can also interpret $R_{n}(H)$ as counting matrices in ${\mathcal{M}}_{n}({\mathbb{Z}};H)$ for which the characteristic polynomial has discriminant equal to zero. We adapt this point of view and count more generally matrices in ${\mathcal{M}}_{n}({\mathbb{Z}};H)$ for which the characteristic polynomial has a given discriminant.
For $A\in{\mathcal{M}}_{n}({\mathbb{Z}};H)$, let $f_{A}\in{\mathbb{Z}}[X]$ be its characteristic polynomial and we define the discriminant of $A$, denoted $\operatorname{discr}(A)$, as the discriminant of $f_{A}$. For $d\in{\mathbb{Z}}$ fixed, we define
${\mathcal{R}}_{n}(d,H)=\{A\in{\mathcal{M}}_{n}({\mathbb{Z}};H):\penalty 10000\ \operatorname{discr}(A)=d\}$
and $R_{n}(d,H)=\#{\mathcal{R}}_{n}(d,H)$. If $d=0$, then $R_{n}(H)=R_{n}(0,H)$.
In this paper we present upper bounds for $R_{n}(d,H)$, significantly improving (1.1) for $d=0$. It appears that for $d\neq 0$ this problem has not been studied before. In this case, our bound is weaker, however stronger than the one of the strength of (1.1) (note that the approach of [19] can be adapted to $d\neq 0$ without any problems).
1.2. Main results
We recall that the notations $U\ll V$ and $V\gg U$ are both equivalent to the statement $|U|\leq cV$, for some constant $c>0$, which throughout this work may depend only on $n$.
Our first main result, using a different approach, gives a significant improvement of (1.1).
Theorem 1.1.
We have
$H^{(n^{2}+n-2)/2}\ll N_{n}(H)\leq R_{n}(H)\ll H^{n^{2}-\Delta_{n}}\log H,$
where
$\Delta_{n}=\max_{r=1,\ldots,n}\min\left\{n-\frac{r(r+1)}{2},r+1\right\},$
while for $n=2,5,7,8$, we also have
$R_{2}(H)\leq H^{2+o(1)},$ $R_{5}(H)\ll H^{5^{2}-4}\log H,$
$R_{7}(H)\ll H^{7^{2}-4}\log H,$ $R_{8}(H)\ll H^{8^{2}-5}\log H.$
We notice that for other small values of $n$ (for which the special treatment of Theorem 1.1 does not apply) we have
$\Delta_{3}=\Delta_{4}=2,\qquad\Delta_{6}=\Delta_{9}=3,\qquad\Delta_{10}=4,$
while for large $n$ we have
$\lim_{n\to\infty}\Delta_{n}/\sqrt{2n}=1.$
We also note that for $n=2$, the lower and upper bounds in Theorem 1.1 agree, that is, one has
$H^{2}\ll N_{2}(H)\leq R_{2}(H)\leq H^{2+o(1)},$
confirming (1.2).
The proof of Theorem 1.1 is based on studying possible factorisations of the characteristic polynomials of matrices and then either using a generalisation of a result from [30] or using modular reduction modulo an appropriately chosen prime $p$. After this we can use known results about the distribution of characteristic polynomials of matrices over finite fields [26].
The above approach, which we use for $d=0$ does not apply for $d\neq 0$. In this case we study the absolute irreducibility of the corresponding variety and use the bound of Browning, Heath-Brown and Salberger [7, Corollary 2] (see also [29]) to prove the following estimate.
Theorem 1.2.
Uniformly over $d\neq 0$, as $H\to\infty$, we have
$R_{n}(d,H)\ll H^{n^{2}-2+o(1)}.$
We note that one of our auxiliary results, an upper bound of the number of matrices with an eigenvalue of degree at most $r$ over $\mathbb{Q}$}, see Lemma 2.5, improves a recent result of Anderson and O’Dorney [3] in some ranges of $r$, which are important for our argument. We also discuss some other applications of our technical tools, in particular improving a result of Assing, Blomer and Nelson [6, Corollary 1.7] in some ranges of parameters, see Section 5.
1.3. Some comments and open questions
We note that our approach does not seem to work for $\operatorname{SL}_{n}({\mathbb{Z}})$ matrices. However, an upper bound
(1.3) ${\mathcal{L}}_{n}\left(d,H\right)\ll H^{n^{2}-n-1/\left(4\left\lceil(n-1)/2\right\rceil\left(n^{2}+n\right)\right)}$
on the cardinality of
${\mathcal{L}}_{n}\left(d,H\right)=\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)\cap\operatorname{SL}_{n}({\mathbb{Z}}):\penalty 10000\ \operatorname{discr}(A)=d\}$
is essentially implicit in [15, Example 1.7] (at least for $d=0$) and is derived in [15, Section 3] from [15, Theorem 3.1]. We recall that by [12, Example 1.6] we have
$H^{n^{2}-n}\ll\#\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)\cap\operatorname{SL}_{n}({\mathbb{Z}})\}\ll H^{n^{2}-n}.$
The only missing ingredient in [15] is a justification of the fact the affine variety defined by
$\det\mathbf{X}=1,\quad\operatorname{discr}(\mathbf{X})=d,\qquad\qquad\mathbf{X}\in{\mathbb{C}}^{n\times n},$
is of dimension $n^{2}-2$. To see that we use the multilinear equation $\det\mathbf{X}=1$ to express $x_{11}$ via the remaining $n^{2}-1$ variables and substitute this in $\operatorname{discr}(\mathbf{X})=d$. This leads to a polynomial equation in $n^{2}-1$ variables. This equation is non-trivial since otherwise all $\operatorname{SL}_{n}({\mathbb{C}})$ matrices have the same discriminant $d$, which is false. Indeed, the identity matrix has a vanishing discriminant, while the diagonal matrix with roots of unity $\exp(2\pi i\nu/n)$, $\nu=1,\ldots,n$, on the main diagonal has a non-zero discriminant.
It is certainly very interesting to investigate whether the methods of [12, 16] apply to estimating $R_{n}(H)$.
Similarly to our conjectured bound (1.2), one can expect that (1.3) can be improved. However before making any conjectures we need to understand the number of monic polynomials $f$ of degree $n$ over $\mathbb{Z}$}, which are similar to those in Remark 2.2 below, but have the constant coefficient $f(0)=1$.
Another interesting direction, is obtaining analogues of our results for some other special classes of matrices, such as symmetric matrices. It seems plausible that our techniques apply to this case as well.
2. Preliminaries on characteristic polynomial, determinants and discrimiants
2.1. Counting matrices with a given characteristic polynomial
For a given polynomial $f\in{\mathbb{Z}}[X]$ let $P_{n}(H;f)$ denote the number of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with a characteristic polynomial $f$.
Some upper bounds on $P_{n}(H;f)$ can be found in [18]. Here we give a short proof in the case of $n=2$, which we need here.
Lemma 2.1.
Uniformly over polynomials $f\in{\mathbb{Z}}[X]$ of degree $\deg f=2$ we have
$P_{2}(H;f)\leq H^{1+o(1)},\qquad H\to\infty.$
Proof.
Let
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$.
If the characteristic polynomial of $A$ is $f(X)=X^{2}-uX+v$, this fixes the trace and the determinant of $A$. Thus we have
$a+d=u\qquad\mbox{and}\qquad ad-bc=v.$
Therefore bc=a(ua)vbc=a(u-a)-v, and when the triple $(a,b,c)$ is fixed, $d$ is uniquely defined.
There are at most $2$ values of $a$ with a(ua)v=0a(u-a)-v=0, in which case either $b=0$ of $c=0$ (or both). So the are at most $4H$ such matrices.
Otherwise for each $a$, the classical bound on the divisor function, see [20, Equation (1.81)], implies that there are at most $H^{o(1)}$ choices for $(b,c)$. So the are at most $H^{1+o(1)}$ such matrices.   $\sqcap\sqcup$
Remark 2.2.
It has been conjectured in [18, Conjecture 1.1] that
(2.1) $P_{n}(H;f)\leq H^{n(n-1)/2+o(1)},$
as $H\to\infty$, uniformly with respect to $f$. Therefore, since $R_{n}(H)$ counts matrices in ${\mathcal{M}}_{n}({\mathbb{Z}};H)$ with a multiple eigenvalue, it is natural to expect that matrices with only one double eigenvalue dominate the count, and thus their characteristic polynomial is of the form $f(X)=(X-a)^{2}g(X)$ where
$a\in{\mathbb{Z}}\quad\text{and}\quad g(X)=X^{n-2}+b_{1}X^{n-3}+\dots+b_{1}X+b_{n-2}\in{\mathbb{Z}}[X],$
with $a\ll H$ and $b_{i}\ll H^{i}$, $i=1,\ldots,n-2$. Therefore, there are $O\left(H^{1+(n-1)(n-2)/2}\right)$ such polynomials $f$, and applying the bound (2.1), we obtain
$R_{n}(H)\leq H^{1+(n-1)(n-2)/2}\cdot H^{n(n-1)/2+o(1)}=H^{n^{2}-2n+2+o(1)}.$
For a prime $p$, we also recall a result of Reiner [26] that gives an explicit formula on the number $P_{n,p}(f)$ of matrices with a given characteristic polynomial over the field ${\mathbb{F}}_{p}$ of $p$ elements. However, for our purposes, we only need the following asymptotical formula for $P_{n,p}(f)$.
Lemma 2.3.
For any monic polynomial $f\in{\mathbb{F}}_{p}[X]$ of degree $n$ we have
$P_{n,p}(f)=\left(1+o(1)\right)p^{n^{2}-n},\qquad\text{as}\ p\to\infty.$
Note that in our argument we only use Lemma 2.3 in the form of an upper bound $P_{n,p}(f)=O\left(p^{n^{2}-n}\right)$.
2.2. Determinants of shifted matrices
The following result follows from a slight extension of the argument of the proof of [30, Theorem 4], which is given for an integral $n\times n$ matrix $K$. However, the proof, which is based on eliminating $K$, never appeals to its integrality. In particular, we have.
Lemma 2.4.
Uniformly over $n\times n$ matrices $K\in{\mathbb{C}}^{n\times n}$ and complex $\alpha\in{\mathbb{C}}$, for
$J_{n}(K;H,\alpha)=\#\left\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right):\penalty 10000\ \det(A-K)=\alpha\right\}$
we have
$J_{n}(K;H,\alpha)\ll H^{n^{2}-n}\log H.$
Proof.
Assume $A,B\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ are obtained from an $n\times(n-1)$-matrix $R$ by augmenting it by two vectors $\mathbf{a},\mathbf{b}\in{\mathbb{Z}}^{n}$, respectively. That is,
$A=\left(R|\mathbf{a}\right)\qquad\text{and}\qquad B=\left(R|\mathbf{b}\right).$
If $\det(A-K)=\det(B-K)$ then putting
$\mathbf{c}=\mathbf{a}-\mathbf{b}\qquad\text{and}\qquad C=\left(R|\mathbf{c}\right),$
we deduce that $\det(C-K_{1})=0$, where the matrix $K_{1}$ is obtained from $K$ by replacing its $n$th column by a zero vector (to see this it is enough to expand $C$ with respect to the last column).
Let
$K=\left(\kappa_{ij}\right)_{i,j=1}^{n}.$
Therefore, for any $R$ and $\alpha$, we have
(2.2) $\begin{split}\#\{\mathbf{a}=(a_{1}&,\ldots,a_{n})\in{\mathbb{Z}}^{n}:\\ &\det\left(\left(R|\mathbf{a}\right)-K\right)=\alpha,\ |a_{i}|<H,\ i=1,\ldots,n\}\\ &\qquad\leq\#\{\mathbf{c}=(z_{1},\ldots,z_{n})\in{\mathbb{Z}}^{n}:\\ &\qquad\qquad\det\left(C-K_{1}\right)=0,\ |c_{i}|<2H,\ i=1,\ldots,n\}.\end{split}$
Indeed, if $\mathbf{a}_{1},\ldots,\mathbf{a}_{J}$ is the list of elements from the first set in (2.2) then the vectors $\mathbf{c}_{j}=\mathbf{a}_{1}-\mathbf{a}_{j}$, $j=1,\ldots,J$, are distinct and belong to the second set in (2.2).
Summing (2.2) over all integral matrices $R=(r_{ij})_{i,j=1}^{n,n-1}$ such that
$|r_{ij}|\leq H,\qquad 1\leq i\leq n,\ 1\leq j\leq n-1,$
and observing that $C\in{\mathcal{M}}_{n}\left({\mathbb{Z}};2H\right)$, we obtain
(2.3) $J_{n}(K;H,\alpha)\leq J_{n}(K_{1};2H,0).$
Repeating the same argument with respect to the (n1)(n-1)th column of the matrix $K_{1}$, we obtain from (2.3)
$J_{n}(K;H,\alpha)\leq J_{n}(K_{2};4H,0),$
where now $K_{2}$ is obtained from $K$ by replacing its (n1)(n-1)th and $n$th columns by a zero vector.
Continuing the same procedure, after $n$ steps we arrive to the inequality
(2.4) $J_{n}(K;H,\alpha)\leq J_{n}(K_{n};2^{n}H,0)$
where $K_{n}=O_{n}$ is the zero matrix. By a result of Katznelson [22] we have
$J_{n}(K_{n};2^{n}H,0)\ll H^{n^{2}-n}\log H,$
which together with (2.4) concludes the proof.   $\sqcap\sqcup$
2.3. Matrices with eigenalues of low degree
Let $R_{n,r}(H)$ be the number of matrices ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ whose characteristic polynomial has an irreducible factor of degree $r\leq n/2$. Anderson and O’Dorney [3] have shown that
$R_{n,r}(H)\ll H^{n^{2}-(n-r)/(n+r-1)}.$
However for our purpose we need the following bound which is better for small values of $r$.
Lemma 2.5.
For any positive integer $r\leq n/2$ we have
$R_{n,r}(H)\ll H^{n^{2}-n+r(r+1)/2}\log H.$
Proof.
By the famous result of Gerschgorin [14], all eigenvalues of a matrix $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ are of order $O(H)$. Hence, if $\lambda$ is an eigenvalue of $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ of degree $d=[{\mathbb{Q}}(\lambda):{\mathbb{Q}}]$ then, since its conjugates are of order $O(H)$, the minimal polynomial $F$ of $\lambda$ over $\mathbb{Q}$},
$F(X)=X^{d}+c_{1}X^{d-1}+\ldots+c_{d-1}X+c_{d}\in{\mathbb{Z}}[X]$
has coefficients satisfying
$c_{i}=O\left(H^{i}\right),\qquad i=1,\ldots,d.$
Thus there are at most $O\left(H^{d(d+1)/2}\right)$ possibilities for $F$.
We now fix some positive integer $r\leq n/2$. We see that the set $\mathcal{L}$} of all possible eigenvalues of degree ${\mathbb{Q}}(\lambda):{\mathbb{Q}}]\leq r$ of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ is of size
(2.5) ${\mathcal{L}}=O\left(H^{r(r+1)/2}\right).$
For each $\lambda\in{\mathcal{L}}$, since
$\det(A-\lambda I_{n})=0,$
where $I_{n}$ is the $n\times n$ identity matrix, we conclude by Lemma 2.4 that there are $O\left(H^{n^{2}-n}\log H\right)$ possible matrices $A$.
Hence, recalling (2.5), we see that the total number of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$, which have an eigenvalue $\lambda$ of degree ${\mathbb{Q}}(\lambda):{\mathbb{Q}}]\leq r$ over $\mathbb{Q}$} is
$R_{n,r}(H)\ll\#{\mathcal{L}}\cdot H^{n^{2}-n}\log H\ll H^{n^{2}-n+r(r+1)/2}\log H,$
which concludes the proof.   $\sqcap\sqcup$
2.4. Absolute irreducibility of shifts of the discriminant of a generic matrix
For a given $d\in{\mathbb{Z}}$, we treat
$\operatorname{discr}(\mathbf{X})-d\in{\mathbb{Z}}\left[\mathbf{X}\right],\qquad\mathbf{X}=\left(X_{ij}\right)_{1\leq i,j\leq n},$
as a polynomial over $\mathbb{Z}$} in $n^{2}$ variables.
We need the following result only for $d\neq 0$, but since it is of independent interest, we prove it in full generality.
Lemma 2.6.
For any $d\in{\mathbb{Z}}$, the polynomial $\operatorname{discr}(\mathbf{X})-d$ is absolutely irreducible over $\mathbb{Q}$}.
Proof.
First we consider $d=0$. Assume that $\operatorname{discr}(\mathbf{X})$ can be factored over the algebraic closure of $\mathbb{Q}$} as
(2.6) $\operatorname{discr}(\mathbf{X})=\prod_{\nu=1}^{s}F_{\nu}(\mathbf{X})^{k_{\nu}},$
into $s$ distinct absolutely irreducible polynomials $F_{1},\ldots,F_{s}\in K[\mathbf{X}]$, defined over some number field $K$, with some multiplicities $k_{1},\ldots,k_{s}\geq 1$. Our goal is to show first that $s=1$ and then that $k_{1}=1$.
We first note that for any integer $m\geq 1$, there exists an integer $a$ such that $\gcd\left(a(2a+1),m\right)=1$. In fact a=m1a=m-1 satisfies this condition.
We now use the above observation with m=n(n1)/2m=n(n-1)/2 and fix some integer $a$ with $\gcd\left(a(2a+1),n(n-1)/2\right)=1$. By the Chebotarev Density Theorem, there are infinitely many primes in the progression $p\equiv 2a+1\pmod{n(n-1)}$ which split completely in the field $K$.
We also choose $p$ to be large enough so that by the Ostrowski’s theorem (see [28, Corollary 2B]), the factorisation (2.6) induces the factorisation
(2.7) $\operatorname{discr}(\mathbf{X})\equiv\prod_{\nu=1}^{s}\overline{F}_{\nu}(\mathbf{X})^{k_{\nu}}\pmod{{\mathcal{P}}},\qquad\overline{F}_{1},\ldots,\overline{F}_{s}\in{\mathbb{F}}_{p}[\mathbf{X}],$
in distinct absolutely irreducible polynomials for any prime ideal $\mathcal{P}$} in $K$ lying over $p$.
We now observe that the factorisation (2.7), together with the celebrated result of Lang and Weil [23] implies that
(2.8) $\{A\in{\mathbb{F}}_{p}^{n\times n}:\penalty 10000\ \operatorname{discr}(A)=0\}=sp^{n^{2}-1}+O\left(p^{n^{2}-3/2}\right).$
On the other hand, by Lemma 2.3 we have
(2.9) $\{A\in{\mathbb{F}}_{p}^{n\times n}:\penalty 10000\ \operatorname{discr}(A)=0\}=(1+o(1))p^{n^{2}-n}N_{0},$
where $N_{0}$ is the number of monic non-squarefree polynomials in ${\mathcal{P}}_{n}({\mathbb{F}}_{p})$ and thus by the classical result of Carlitz [10, Section 6], for any $n\geq 2$, we have
(2.10) $N_{0}=p^{n-1}.$
Combining (2.8), (2.9) and (2.10) and taking $p$ to be sufficiently large we see that $s=1$. Hence, we can write (2.6) as
$\operatorname{discr}(\mathbf{X})=F(\mathbf{X})^{k},$
with an absolutely irreducible polynomial $F\in K[\mathbf{X}]$.
Finally, we want to conclude that $k=1$, that is, $\operatorname{discr}(\mathbf{X})$ is absolutely irreducible. It is enough to show that $\operatorname{discr}(\mathbf{X})$ cannot be a power of a polynomial by considering some specialisation of the variables $\mathbf{X}$.
In particular, we consider the matrix
$\mathbf{X}^{*}={\mathrm{diag}}(J,2,\ldots,n-1),$
where
$J=\left(\begin{array}[]{ll}0&X_{1,2}\\ X_{2,1}&0\end{array}\right).$
From the Laplace expansion with respect to the first two rows, we see that the characteristic polynomial of $\mathbf{X}^{*}$ is
$f(Y)=\det(\mathbf{X}^{*}-Y\cdot I_{n})=g(Y)h(Y)\in R[Y]$
where $R=K[X_{1,2},X_{2,1}]$ and
$g(Y)=\det(D-Y\cdot I_{2})=Y^{2}-X_{1,2}X_{2,1},$
$h(Y)=\det{\mathrm{diag}}(2-Y,\ldots,n-1-Y)=\prod_{i=2}^{n-1}\left(i-Y\right).$
The discriminants of $g$ and $h$ are
$\operatorname{discr}(g)=4X_{1,2}X_{2,1}\qquad\mbox{and}\qquad\operatorname{discr}(h)=\prod_{2\leq i<j\leq n-1}(i-j)^{2},$
and the resultant of $g$ and $h$ is given by
$\mathrm{Res}(g,h)=\prod_{i=2}^{n-1}\left(i^{2}-X_{1,2}X_{2,1}\right).$
Recall that
$\operatorname{discr}(f)=\operatorname{discr}(g)\operatorname{discr}(h)\mathrm{Res}(g,h)^{2}.$
Since $\operatorname{discr}(g)$ and $\mathrm{Res}(g,h)$ are relatively prime polynomials in $R$, and $\operatorname{discr}(g)$ is squarefree while $\operatorname{discr}(h)$ is a scalar in $R$, we deduce that $\operatorname{discr}(f)$ cannot be a power of a polynomial in $R$. Hence for $d=0$ we have $k=s=1$.
We now consider $d\neq 0$. Assume that
$\operatorname{discr}(\mathbf{X})-d=GH,$
where $G,H\in{\mathbb{F}}_{q}[\mathbf{X}]$ are of degrees $k,\ell\geq 1$, respectively. Since $\operatorname{discr}(\mathbf{X})$ is a homogeneous polynomial, we obtain that
$\operatorname{discr}(\mathbf{X})=G_{k}H_{\ell},$
where $G_{k}$ and $H_{\ell}$ are the top degree $k$ and $\ell$ homogeneous parts of $G$ and $H$, respectively. However, this contradicts the already established absolute irreducibility of $\operatorname{discr}(\mathbf{X})$, which completes the proof.   $\sqcap\sqcup$
3. Proof of Theorem 1.1
3.1. The general upper bound
If $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ has all eigenvalues of degree at least $r+1$ and at least one multiple eigenvalue then its characteristic polynomial $f\in{\mathbb{Z}}[X]$ is of the form $f=g^{2}h$ with monic polynomials
$g,h\in{\mathbb{Z}}[X]\qquad\mbox{and}\qquad\deg g\geq r+1.$
We now choose $p$ as the smallest prime with $p\geq 2H+1$. Clearly the reduction of $A$ modulo $p$ has characteristic polynomial $f_{p}\in{\mathbb{F}}_{p}[X]$ is of the form $f_{p}=g_{p}^{2}h_{p}$ with monic polynomials
$g_{p},h_{p}\in{\mathbb{Z}}[X]\qquad\mbox{and}\qquad\deg g_{p}\geq r+1.$
Thus there are
$\sum_{r+1\leq d\leq n/2}p^{d}\cdot p^{n-2d}=\sum_{r+1\leq d\leq n/2}p^{n-d}\leq 2p^{n-r-1}$
possibilities of $f_{p}$. By Lemma 2.3 the number of $n\times n$ matrices over ${\mathbb{F}}_{p}$ with characteristic polynomials of this type is
(3.1) $O\left(p^{n^{2}-n}\cdot p^{n-r-1}\right)=O\left(p^{n^{2}-r-1}\right).$
Since $p\geq 2H+1$, each of these matrices over ${\mathbb{F}}_{p}$ can be lifted to at most one matrix $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$. Furthermore, since by our choice, $p=O(H)$, we see from (3.1) that there are
(3.2) $Q_{n,r}(H)=O\left(H^{n^{2}-r-1}\right)$
such matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$.
Combining Lemma 2.5 and (3.2), we derive
(3.3) $R_{n}(H)\leq R_{n,r}(H)+Q_{n,r}(H)\ll H^{n^{2}-n+r(r+1)/2}\log H+H^{n^{2}-r-1}.$
Since $r$ is arbitrary, the result follows.
3.2. The cases $n=2,5,7,8$
When $n=2$, then any matrix counted by $R_{2}(H)$ with eigenvalue $\lambda$ has the characteristic polynomial of the form
$(X-\lambda)^{2}.$
Since there are $O(H)$ possibilities for $\lambda$, applying Lemma 2.1, we conclude that for each possible eigenvalue $\lambda$ there are $H^{1+o(1)}$ matrices with characteristic polynomial $(X-\lambda)^{2}$. Therefore,
$R_{2}(H)\leq H\cdot H^{1+o(1)}=H^{2+o(1)}.$
The improvement for the cases $n=5,7,8$ comes from the situation when $r$ satisfies the inequalities
(3.4) $r+1>n/3\qquad\mbox{and}\qquad(r+1)(r+2)<2n.$
This enforces $n<15$.
Indeed, following the proof in Section 3.1, we note that, in the case when all eigenvalues are of degree at least $r+1$ and at least one is a multiple eigenvalue, then its characteristic polynomial $f\in{\mathbb{Z}}[X]$ is of the form $f=g^{2}h$ with monic polynomials
$g,h\in{\mathbb{Z}}[X]\qquad\mbox{and}\qquad\deg g,\deg h\geq r+1.$
Looking at the degree, one needs $n\geq 2(r+1)+r+1$, that is, $r+1\leq n/3$. Therefore, if $r+1>n/3$ such a factorisation of $f$ cannot occur. Hence for such pairs $(n,r)$ we simply have
$R_{n}(H)\leq R_{n,r}(H)\ll H^{n^{2}-n+r(r+1)/2}\log H.$
The second inequality in (3.4) comes from imposing that the bound obtained on $R_{n,r}(H)$ in Lemma 2.5 is smaller than the bound (3.2) on $Q_{n,r}(H)$ and hence the above bound is stronger than (3.3),
Simple calculations show that this is possible for $n=5$ (with $r=1$), $n=7$ (with $r=2$) and $n=8$ (with $r=2$), and we get the desired upper bounds on $R_{5}(H)$, $R_{7}(H)$ and $R_{8}(H)$.
3.3. The lower bound
The lower bound comes from the following simple construction: we consider $n\times n$ triangular matrices
$A=\begin{pmatrix}a&b&*&\cdots&*\\ 0&a&*&\cdots&*\\ 0&0&a_{3,3}&\cdots&*\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ 0&0&0&\cdots&a_{n,n}\end{pmatrix}$,
with $a,b\in{\mathbb{Z}}$, $b\neq 0$, $|a|,|b|\leq H$ and $a_{i,i}\neq a$, $|a_{i,i}|\leq H$, $i=3,\ldots,n$.
Clearly, the minor $M_{2,1}$ of
$A-a\cdot I_{n}=\begin{pmatrix}0&b&*&\cdots&*\\ 0&0&*&\cdots&*\\ 0&0&a_{3,3}-a&\cdots&*\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ 0&0&0&\cdots&a_{n,n}-a\end{pmatrix}$,
(that is, the submatrix of $A-a\cdot I_{n}$ obtained by removing the second row and the first column) is an upper triangular matrix with non-zero elements on the main diagonal, and hence it is non-singular. Therefore, $A-a\cdot I_{n}$ is of rank n1n-1 and thus the geometric multiplicity of the eigenvalue $a$, satisfies
$\dim{\mathrm{Ker}}(A-a\cdot I_{n})=1.$
Since this is different from the algebraic multiplicity of $a$, which is $2$ (the multiplicity of $a$ as a root of the characteristic polynomial), the matrix $A$ is not diagonalisable.
Therefore,
$R_{n}(H)\geq N_{n}(H)\gg H\cdot H^{n-2}\cdot H^{n(n-1)/2}=H^{(n^{2}+n-2)/2},$
which concludes the proof.
4. Proof of Theorem 1.2
4.1. The cases $n=2$
The characteristic polynomial of
$A=\begin{pmatrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{pmatrix}$,
is $X^{2}-\left(a_{11}+a_{22}\right)X+\left(a_{11}a_{22}-a_{12}a_{21}\right)$ and hence its discriminant is
$\left(a_{11}+a_{22}\right)^{2}-4\left(a_{11}a_{22}-a_{12}a_{21}\right)=\left(a_{11}-a_{22}\right)^{2}+4a_{12}a_{21}.$
Considering the cases $\left(a_{11}-a_{22}\right)^{2}=d$ and $\left(a_{11}-a_{22}\right)^{2}\neq d$ separately, and using the classical bound on the divisor function, see [20, Equation (1.81)], we conclude the proof in this case.
4.2. Integer points on varieties
We recall the following general bound of Salberger [29, Theorem 7.4] (see also Browning, Heath-Brown and Salberger [7, Corollary 2 and Lemma 8]), which in turn improves the result of Pila [25, Theorem A] under some additional constraints.
Lemma 4.1.
Let $F(X_{1},\ldots,X_{m})\in{\mathbb{Z}}[X_{1},\ldots,X_{m}]$ be a polynomial of degree $\deg F\geq 4$ in $m\geq 3$ variables and with an absolutely irreducible highest form. Then
$\#\{(x_{1},\ldots,x_{m})\leq\left({\mathbb{Z}}\cap[-H,H]\right)^{m}:\penalty 10000\ F(x_{1},\ldots,x_{m})$ $=0\}$
$\leq H^{m-2+o(1)},$
as $H\to\infty$.
4.3. Concluding the proof
Let now $n\geq 3$. We observe that the characteristic polynomial of a matrix
$\mathbf{X}=\left(X_{ij}\right)_{1\leq i,j\leq n}$
whose entries are considered as variables, is of the form
$f_{\mathbf{X}}(Z)=Z^{n}+F_{1}Z^{n-1}+\ldots+F_{n-1}Z+F_{n},$
where $F_{i}\in{\mathbb{Z}}[\mathbf{X}]$ is a polynomial in $n^{2}$ variables of total degree $i$, $i=1,\ldots,n$. Hence, $\operatorname{discr}\left(f_{\mathbf{X}}\right)$ is also a polynomial in $X_{ij}$, $1\leq i,j\leq n$.
We now investigate the degree of $\operatorname{discr}\left(f_{\mathbf{X}}\right)$ as a polynomial in $X_{ij}$, $1\leq i,j\leq n$. Note that in order to apply Lemma 4.1 we only need a lower bound on $\deg\operatorname{discr}\left(f_{\mathbf{X}}\right)$. To obtain such a bound we consider a symmetric matrix
$\widetilde{\mathbf{X}}=\left(\widetilde{X}_{ij}\right)_{1\leq i,j\leq n},\qquad\widetilde{X}_{ij}=\widetilde{X}_{ji},\ 1\leq i,j\leq n,$
and note that
[…]
On the other hand, it has been shown by Lax [24] that
(4.2) $\deg\operatorname{discr}\left(f_{\widetilde{\mathbf{X}}}\right)=n(n-1).$
Since $n\geq 3$, by Lemma 2.6 and also using (4.1) and (4.2), we can apply Lemma 4.1 with $F$ replaced by $\operatorname{discr}\left(f_{\mathbf{X}}\right)-d$ and $m=n^{2}$, which instantly implies the desired result.
Remark 4.2.
While we are only interested in a lower bound on the degree of the discrimiant $\deg\operatorname{discr}\left(f_{\mathbf{X}}\right)$, we expect that (4.1) is actually an equality and thus by (4.2) we have $\deg\operatorname{discr}\left(f_{\mathbf{X}}\right)=n(n-1)$. The exact evaluation of $\deg\operatorname{discr}\left(f_{\mathbf{X}}\right)$ is a question of independent interest.
5. Other applications of our approach
5.1. Matrix lifting
We note that the generalisation of the result from [30], given in Lemma 2.4, allows us to study modular lifts of matrices with a prescribed determinant. Namely, given an integer $q$ and a matrix $B=\left(b_{ij}\right)_{1\leq i,j\leq n}$ with $b_{ij}\in\{0,\ldots,q-1\}$, $1\leq i,j\leq n$, we consider the set
${\mathcal{T}}_{n}(d,H;q,B)=\{A\in{\mathcal{M}}_{n}({\mathbb{Z}};H):\penalty 10000\ A\equiv B\pmod{q},\ \det A=d\}.$
For $d=1$, that is for $\operatorname{SL}_{n}({\mathbb{Z}})$ matrices, the question about the smallest $H$ (as a function of $q$) for which ${\mathcal{T}}_{n}(1,H;q,B)\neq\emptyset$ for any $B$ has recently attracted a lot of attention, see [5, 21] and references therein. This can also be asked for other values of $d$. Here we remark that Lemma 2.4 applied with $K=q^{-1}B$ and $\alpha=d/q^{n}$ (and with $\left\lceil H/q\right\rceil$ instead of $H$) immediately implies that for $H\geq 2q$ we have
(5.1) ${\mathcal{T}}_{n}(d,H;q,B)\ll(H/q)^{n^{2}-n}\log(H/q)$
uniformly over $d$ and $B$. This can be compared with the result of Assing, Blomer and Nelson [6, Corollary 1.7] (see also [5, Theorem 1.4]) which covers the case when $d=1$ and $B=I_{n}$ is equivalent to the bound
(5.2) ${\mathcal{T}}_{n}(1,H;q,I_{n})\leq\left(\frac{H^{n^{2}-n}}{q^{n^{2}-1}}+H^{(n^{2}-n)/2}\right)H^{o(1)}.$
The bound (5.1) is more general and stronger than (5.2) for large $q$, namely, for
$q\geq H^{1/2+\varepsilon}$
for any fixed $\varepsilon>0$.
5.2. Singular matrix polynomials
We now give an applications of Lemma 2.4 to counting singular matrix polynomials $F(A)$ with a given $F\in{\mathbb{C}}[X]$ and $A\in{\mathcal{M}}_{n}({\mathbb{Z}};H)$. More precisely, we have
(5.3) $\{A\in{\mathcal{M}}_{n}({\mathbb{Z}};H):\penalty 10000\ \det F(A)=0\}\leq H^{n^{2}-n+o(1)}.$
Indeed, the eigenvalues of $F(A)$ are of the shape $F(\lambda)$, where $\lambda$ runs through the eigenvalues of $A$. Hence, if $\det F(A)=0$ then we have $\det(A-\rho I_{n})=0$ where $\rho$ is one of the roots of $F$ and as before, $I_{n}$ denotes the $n\times n$ identity matrix. Invoking Lemma 2.4, we obtain (5.3).
5.3. Condition number of integral matrices
We recal that the condition number of an $n\times n$ matrix $A$ (with arbitrary complex entries) is given by
$\kappa(A)=\frac{\max_{i=1,\ldots,n}\sigma_{i}(A)}{\min_{i=1,\ldots,n}\sigma_{i}(A)},$
where $\sigma_{1}(A),\ldots,\sigma_{n}(A)$ are the singular values of $A$, that is,
$\sigma_{i}(A)=\sqrt{\lambda_{i}\left(AA^{T}\right)},$
where $\lambda_{i}\left(AA^{T}\right)$ are the eigenvalues of $AA^{T}$, $i=1,\ldots,n$, and $A^{T}$ denotes the transposition of $A$. The condition number is responsible for the numerical stability of many standard linear algebra algorithms (the smaller the better) and its distribution has been studied for a large variety of families of matrices, see [4, 9, 11, 17, 27, 31, 32, 33, 34] and references therein. Here we address a discrete analogue of this question. For example, one can think about random matrices from ${\mathcal{M}}_{n}({\mathbb{Z}};H)$ as obtained from random real matrices after decimal approximations of their entries, with a given precision, and then scaling the new entries to make them integer.
Since for $A\in{\mathcal{M}}_{n}({\mathbb{Z}};H)$, the entries of $AA^{T}$ are of size $O(H^{2})$, by [14] we conclude that
$\max_{i=1,\ldots,n}\sigma_{i}(A)\ll H.$
This also implies that if $\min_{i=1,\ldots,n}\sigma_{i}(A)\leq H/L$ then
$|\det A|=\prod_{i=1}^{n}\sigma_{i}(A)\ll L^{-1}H^{n}.$
Thus, we see from Lemma 2.4, or [30, Theorem 4], that for all but $O\left(L^{-1}H^{n^{2}}\log H\right)$ matrices $A\in{\mathcal{M}}_{n}({\mathbb{Z}};H)$, we have $\kappa(A)\leq L$. For example, taking $L=\log H\log\log H$, we see that for almost all matrices from ${\mathcal{M}}_{n}({\mathbb{Z}};H)$ the condition number is reasonably small, namely, at most $\log H\log\log H$.
Acknowledgement
The authors would like to thank Wadim Zudilin for useful comments and encouragement.
This work was supported, in part, by the Australian Research Council Grants DP230100530 and DP230100534.